# Soft Masked Transformer for Point Cloud Processing with Skip Attention-Based Upsampling

### v1

@@ Abstract
Point cloud processing methods leverage local and global point features to cater to downstream tasks, yet they often overlook the task-level context inherent in point clouds during the encoding stage. We argue that integrating task-level information into the encoding stage significantly enhances performance. To that end, we propose SMTransformer which incorporates task-level information into a vector-based transformer by utilizing a soft mask generated from task-level queries and keys to learn the attention weights. Additionally, to facilitate effective communication between features from the encoding and decoding layers in high-level tasks such as segmentation, we introduce a skip-attention-based up-sampling block. This block dynamically fuses features from various resolution points across the encoding and decoding layers. To mitigate the increase in network parameters and training time resulting from the complexity of the aforementioned blocks, we propose a novel shared position encoding strategy. This strategy allows various transformer blocks to share the same position information over the same resolution points, thereby reducing network parameters and training time without compromising accuracy. Experimental comparisons with existing methods on multiple datasets demonstrate the efficacy of SMTransformer and skip-attention-based up-sampling for point cloud processing tasks, including semantic segmentation and classification. In particular, we achieve state-of-the-art semantic segmentation results of 73.4% mIoU on S3DIS Area 5 and 62.4% mIoU on SWAN dataset.
@@
Deep learning, 3D point clouds, Soft mask, Transformer, Attention-based up-sampling.
@@ I Introduction
With the rapid development of 3D sensors such as LiDARs and depth cameras, the accessibility of 3D point cloud data has dramatically increased. Consequently, its importance is growing across various applications, such as autonomous driving [1], [2], robotics[3], and industrial automation[4]. Techniques for processing 3D point clouds have attracted the interest of many researchers. Effectively learning features from 3D point clouds is challenging due to its irregular nature, as opposed to regular grid like structure of images. Inspired by the success of grid convolutions, several methods have been proposed to transform irregular point clouds into regular representations, including projected images[5],[6], or voxels[7], [8]. However, the discretisation process inevitably sacrifices significant geometric information.
@@
Early approaches learn features from raw point clouds by employing a shared multi-layer perceptron (MLP) on each point and use symmetric functions (e.g., max-pooling) to aggregate the most prominent features over the receptive field. However, this design ignores local structures crucial for shape representation. Inspired by 2D grid convolutions, point convolutions utilize correlation functions to quantify connections among points, enabling the model to leverage local features (e.g., edges, corners, surfaces) and contextual information (e.g., scene layout). Some point convolutions use weight functions to learn weights from various local point geometric connections, such as point coordinates, coordinate differences, distances, etc., [9], [10], [11], [12]. Others associate coefficients (derived from point coordinates) [13], [14], [15] with weight functions to adjust the learned weights.
@@
In contrast to point convolution, the point transformer focuses on learning feature connections and attention maps between point features through a scalar or vector-based attention mechanism globally [16], [17], [18], [19], [20], [21] or over a local receptive field [22], [23]. Recently, with the help of position encoding, the aggregation ability of point transformers has seen significantly improvement. While point transformers are powerful learners, they generally fall short of simultaneously exploiting local features and global context, failing to establish communication between the two. Current point convolution and transformer based approaches strive to enhance the encoder and decoder designs to facilitate proficient point connection learning for improved performance on downstream tasks such as segmentation, detection and classification. However, these methods often overlook capturing the task-level contextual information of the entire point cloud which results in sub-optimal performance on the target task, such as inaccurate predictions when segmenting small objects with intricate boundaries.
@@
To address the aforementioned issues, we propose a Soft Masked Transformer (SMTransformer) that incorporates task-level contextual information into attention mechanisms and utilises this information to guide local feature learning softly. Specifically, the SMTransformer predicts task score keys and queries over the global receptive field and then obtains a soft mask from these keys and queries to re-weight the local attention map.
@@
Moreover, in conventional networks, different point transformer blocks often employ distinct position encoding information for the same resolution. This increases the network parameters and the training time. Intuitively, the same resolution points in the network have the same position information, a characteristic independent of the number and location of point transformers.
Motivation and Contributions: The motivation for this paper lies in addressing the limitations of existing point cloud processing methods. While existing methods effectively utilize local and global point features, they often neglect the inherent task-level context in point clouds during the encoding stage. This oversight can lead to sub-optimal performance in downstream tasks. To overcome this challenge, the paper proposes the SMTransformer, which integrates task-level information into the encoding stage by introducing a soft mask generated from task-level queries and keys to adjust the attention weights. Additionally, this paper introduces a skip-attention-based up-sampling block to enhance communication between features from encoding layers over different resolutions, particularly in high-level tasks like segmentation. To further improve efficiency, a novel shared position encoding strategy is proposed to reduce network parameters and training time without sacrificing accuracy. These innovations aim to enhance the performance, effectiveness and efficiency of point cloud processing methods in general. To summarize, our contributions are threefold:
  • We propose a Soft Masked Transformer block, which integrates task-level information into the attention mechanism, enhancing its effectiveness for downstream tasks such as semantic segmentation and classification.
@@
  • We present a shared position encoding strategy, which reduces network parameters by 24.3% and training time by 33.3%. This strategy enhances the efficiency of the network without sacrificing performance.
We conduct extensive experiments on benchmark datasets to showcase the effectiveness of our proposed method and their robust generalization across various tasks, including indoor semantic segmentation, outdoor semantic segmentation, and object classification. Our method consistently achieves competitive results compared to existing point transformer-based approaches. Particularly noteworthy is our method’s achievement of state-of-the-art semantic segmentation performance, attaining the remarkable mIoU of 73.4% on the S3DIS Area 5 and 62.4% on the SWAN dataset without any pre-training.
@@ II Related Work
II-A Point-based Methods
Aiming to maximize the preservation of geometric information in point clouds, state-of-the-art methods prefer to directly process the raw point clouds. The development of the most important unit (i.e., local aggregation) in the point cloud processing network can be broadly divided into three categories explained below.
1) MLP-Based Approaches: PointNet [24] is considered a milestone in point cloud based deep learning. It employs shared MLPs to leverage point-wise features and utilizes a symmetric function such as max-pooling to aggregate these features into global representations. However, the network’s performance is limited as it does not account for the spatial relationships among local points, which are crucial for vision tasks. To address this issue, hierarchical architectures have been proposed to aggregate local features with MLPs [25, 26] such that the model can benefit from efficient sampling and grouping of the point set. Recent works [27, 28, 29] have focused on enhancing point-wise features by hand-crafting geometric connections such as curves, triangles, umbrella orientation, or affine transformations. Additionally, graphs have been introduced to points [30, 20, 31, 32, 33, 34, 35] and subsequent geometric representations, such as edges, contours, curvature, and connectivity. However, these strategies might lack generality due to the need to optimize hyperparameters for hand-crafted representations or graphs across datasets with varying densities or shape styles.
2) Convolution-Based Approaches: Inspired by the success of 2D convolution, various works have successively proposed novel point convolutions on points or point graphs. These methods dynamically learn convolutional weights through functions derived from local geometric connections. A popular category of methods in this domain focuses on designing weight functions. Among these, some approximate the weight function using MLPs [36, 37, 38, 13, 14, 33, 39], spline functions [11], a family of polynomial functions [12], or standard unparameterized Fourier functions [40]. Unlike these methods with dynamic convolution kernels, KCNet [9] and KPConv [10] predefine a set of fixed kernels (i.e., template points) in the local receptive field and then learn the weights on these kernels from the geometric connections between local points and template points using Gaussian and linear correlation functions, respectively. However, the number and position of kernel points need to be optimized for different datasets. Another approach [13, 14, 15] associates coefficients with kernels to further adjust the learned weights, where the coefficients are obtained through kernel density estimation, inverse density functions, and fuzzy functions of point coordinates.
@@ II-A Point-based Methods
3) Transformer-Based Approaches: Unlike convolution-based methods, which learn convolutional weights from low-level point coordinates, attention mechanisms learn attention weights from the connections between point features, thereby exploiting high-level contextual information. Motivated by the success of attention mechanisms in natural language processing and image processing tasks, early methods applied self-attention to global points through scalar dot-product [16, 17, 18, 19, 20, 21], but suffered from high computational costs. These early attention-based methods did not demonstrate superior performance due to the lack of employing position encoding.
The Point Transformer [23] introduces local vector attention to the local points. Additionally, it emphasizes the significance of point position encoding. Subsequent work, Point Transformer V2 [22], incorporates multi-grouping into vector attention inspired by the multi-head strategy. It also enhances position encoding by introducing an additional multiplier to the relation vector, which facilitates learning complex positional relations.
@@
To exploit long-range contextual information, the Stratified Transformer [41] densely selects nearby points over a cubic window and sparsely selects distant points. This stratified strategy enlarges the effective receptive field without incurring too much computational overhead. However, by using a window-based approach, the Stratified Transformer focuses on an expanded local region rather than the global region. Moreover, hyper-parameters such as the window size and number of distant points must be optimized for different datasets with varying densities.
@@ II-B Up-sampling in 3D Point Clouds
The hierarchical architecture of a network is instrumental in learning long-range contextual information through down-sampling and grouping operations. Up-sampling involves interpolating new points between known points and adjusting the features of these new points based on their mutual distance to propagate the learned context features to each point. Compared to the extensive research on down-sampling [14, 42, 21, 18, 43] and grouping operations [25, 26, 20, 44, 45, 46, 34, 33, 37], few works specifically emphasize the up-sampling operation. In the Point Transformer [23], the transition up module uses an interpolation operation to recover new point features from known point features through indexing and then integrates these features with those from the encoding layer via a skip connection. Similarly, in Point Transformer V2 [22], the fusion of encoding and decoding layer features is achieved through a skip connection. The distinguishing factor here is that the new point features are unpooled by grid unpooling instead of interpolation. While these methods are simple, they lack semantic awareness and ignore the contextual connection between the encoding and decoding layers.
Our survey highlights several gaps of in current point cloud processing methods. Their fundamental unit, such as point MLP, convolution, transformer, completely ignores task-level information which then leads to sub-optimal performance in downstream tasks. Moreover, the upsampling methods do not effectively communicate between the encoding layers over different resolution points to refine the context information for high-level tasks. Motivated by these gaps, we propose a soft-masked transformer and skip-attention-based upsampling. Moreover, we propose shared position encoding to reduce the network parameters and training time.
@@ III Method
We revisit the two classical local vector attention-based point transformers, namely Point Transformer and Point Transformer V2, in Section III-A. Then, we present the SMTransformer block in Section III-B, followed by our skip attention-based up-sampling block in Section III-C. We introduce the shared position encoding strategy in Section III-D, and finally, in Section III-E, we provide details about our network.
@@ III-A Rethinking Vector Attention based Transformer
$\mathcal{G}_{i}=\sum_{j=1}^{K}\mathcal{A}\big{(}(K_{ij}^{f}\ominus Q_{i}^{f})\oplus\delta_{b}\big{)}\odot\big{(}V_{ij}^{f}\oplus\delta_{b}\big{)},$ (1)
where $Q_{i}^{f}\in\mathbb{R}^{{N\times C}}$ is the query matrices of $f_{i}$. $K_{ij}^{f}$ and $V_{ij}^{f}\in\mathbb{R}^{{N\times K\times C}}$ are the key and value matrices of $f_{ij}$. The subtraction operation $\ominus$ between $Q_{i}^{f}$ and $K_{ij}^{f}$ is performed via broadcasting to ensure dimension compatibility. $\delta_{b}=\delta_{b}(\Delta p_{ij})$ is the position encoding bias function. $\mathcal{A}(\cdot)$ donates the local vector attention function, implemented by MLP, followed by the softmax. $\sum$ stands for the symmetric operation (SOP) (e.g. summation) and $\odot$ is the element-wise multiplication operation. $\oplus$ is the element-wise addition operation and $\mathcal{G}_{i}$ is the output feature of transformer.
@@
$\mathcal{G}_{i}=\sum_{j=1}^{K}\mathcal{A}\big{(}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{m}}(K_{ij}^{f}\ominus Q_{i}^{f})\oplus\delta_{b}\big{)}\odot\big{(}V_{ij}^{f}\oplus\delta_{b}\big{)},$ (2)
where $\delta_{m}=\delta_{m}(\Delta p_{ij})$ is the position encoding multiplier function. The diagram of the above two classical point transformer are illustrated as Fig. 1.(a)(b). The attention function $\mathcal{A}(\cdot)$ learns robust weights from the rich relationships, including the low-level geometric relationship (i.e. relation position) and the high-level contextual relationship (i.e. relation feature). Despite the overarching influence of the task on the entire network, the fundamental unit transformer overlooks task-level information, resulting in the loss of crucial task-related details (e.g. semantic information) during the encoding and decoding stages.
@@ III-B Soft Masked Transformer Block (SMTB)
We propose a novel soft-masked transformer for point cloud processing. The soft-masked transformer can be expressed as,
@@ III-B1 Soft Mask
where the $W_{q}^{s}$ and $W_{k}^{s}$ ($\mathbb{R}^{N\times C}\rightarrow\mathbb{R}^{N\times T}$) are prediction functions implemented as linear layers followed by softmax, where $T$ is the number of the task classes. $G[\cdot]$ is the grouping operation to obtain the task scores of neighbour points at point $p_{i}$. The soft mask is generated from the score difference $(K_{ij}^{s}-Q_{i}^{s})\in\mathbb{R}^{{N\times K\times T}}$ as,
$S(\cdot)=S(Q_{i}^{s},K_{ij}^{s})=||{\rm Max}({\rm Norm}(K_{ij}^{s}\ominus Q_{i}^{s}))||_{2},$ (5)
where the ${\rm Max}(\cdot)$ is the maximum, ${\rm Norm}(\cdot)$ is Min-Max Normalization, and $||\cdot||_{2}$ is the Euclidean Norm. The soft mask consists of real numbers ranging between 0 and 1, typically representing the probability values of identical labels among different points. This allows the model to assign higher importance to neighbouring points with distinct predicted labels. Taking segmentation as an example, the soft mask enhances the robustness of attention weights around class boundaries. Unlike traditional hard masks (i.e., binary masks), soft masks are more flexible and efficient as they do not require explicit rules or conditions for determination.

### v2

@@ Abstract
Point cloud processing methods leverage local and global point features to cater to downstream tasks, yet they often overlook the task-level context inherent in point clouds during the encoding stage. We argue that integrating task-level information into the encoding stage significantly enhances performance. To that end, we propose SMTransformer which incorporates task-level information into a vector-based transformer by utilizing a soft mask generated from task-level queries and keys to learn the attention weights. Additionally, to facilitate effective communication between features from the encoding and decoding layers in high-level tasks such as segmentation, we introduce a skip-attention-based up-sampling block. This block dynamically fuses features from various resolution points across the encoding and decoding layers. To mitigate the increase in network parameters and training time resulting from the complexity of the aforementioned blocks, we propose a novel shared point position encoding strategy. This strategy allows various transformer blocks to share the same position information over the same resolution points, thereby reducing network parameters and training time without compromising accuracy. Experimental comparisons with existing methods on multiple datasets demonstrate the efficacy of SMTransformer and skip-attention-based up-sampling for semantic segmentation task. In particular, we achieve state-of-the-art semantic segmentation results of 73.9% mIoU on S3DIS Area 5 and 62.4% mIoU on SWAN dataset.
@@
Deep learning, 3D point clouds, Soft mask, Transformer, Attention-based up-sampling. Note to Practitioners—Point cloud processing underpins automation tasks such as robotic perception, navigation, and inspection, where accurate 3D understanding is essential. Existing methods often prioritize vision benchmarks while overlooking automation needs like efficiency on limited hardware and robustness in real-world environments. The proposed SMTransformer embeds task-level guidance into feature learning and employs skip-attention up-sampling to improve segmentation accuracy with practical efficiency. It is well-suited for robotic manipulation, autonomous driving, and inspection applications. Current limitations include reliance on GPUs and sensitivity to extreme density variations. Future work will target edge-device deployment and multi-task extensions.
@@ I Introduction
With the rapid development of 3D sensors such as LiDARs and depth cameras, the accessibility of 3D point cloud data has dramatically increased. Consequently, its role in automation-oriented applications has expanded to areas such as autonomous driving [13], robotics[31], and industrial automation[27], where reliable 3D perception is critical for decision-making and control. Techniques for processing 3D point clouds have attracted the interest of many researchers. Effectively learning features from 3D point clouds is challenging due to its irregular nature, as opposed to regular grid like structure of images. Inspired by the success of grid convolutions, several methods have been proposed to transform irregular point clouds into regular representations, including projected images, or voxels. However, the discretization process inevitably sacrifices significant geometric information.
@@
Early approaches learn features from raw point clouds by employing a shared multi-layer perceptron (MLP) on each point and use symmetric functions (e.g., max-pooling) to aggregate the most prominent features over the receptive field. However, this design ignores local structures crucial for shape representation. Inspired by 2D grid convolutions, point convolutions employ correlation functions to quantify relationships among points, enabling the model to leverage both local features (e.g., edges, corners, surfaces) for robotic perception and contextual cues (e.g., scene layout) essential in automation tasks. Some point convolutions use weight functions to learn weights from various local point geometric connections, such as point coordinates, coordinate differences, distances, etc., [28]. Others associate coefficients (derived from point coordinates) [34] with weight functions to adjust the learned weights.
@@
In contrast to point convolution, the point transformer focuses on learning feature connections and attention maps between point features through a scalar or vector-based attention mechanism globally [7], [40] or over a local receptive field [36]. Recently, with the help of position encoding, the aggregation ability of point transformers has seen significantly improvement. While point transformers are powerful learners, they generally fall short of simultaneously exploiting local features and global context, failing to establish communication between the two. Current point convolution and transformer based approaches strive to enhance the encoder and decoder designs to facilitate proficient point connection learning for improved performance on downstream tasks such as segmentation and classification. However, these methods often overlook capturing the task-level information of the entire point cloud. For instance, task-level information in the semantic segmentation task refers to the semantic priors (e.g. class scores) provided by the complete set of class labels, which are crucial for disambiguating locally similar but semantically different structures
@@
To address the aforementioned issues, we propose a Soft Masked Transformer (SMTransformer) that incorporates task-level contextual information into attention mechanisms and utilizes this information to guide local feature learning softly. Specifically, the SMTransformer predicts class score keys and queries over the global receptive field and then obtains a soft mask from these keys and queries to re-weight the local attention map.
@@
Moreover, in conventional networks, different point transformer blocks often employ distinct position encoding information for the same resolution. This increases the network parameters and the training time. Intuitively, the same resolution points in the network have the same position information, a characteristic independent of the number and location of point transformers. To summarize, our contributions are threefold:
  • We propose a Soft Masked Transformer block, which integrates task-level information into the attention mechanism, enhancing its effectiveness for semantic segmentation.
@@
  • We present a shared point position encoding strategy, which reduces network parameters by 24% and training time by 33%. This strategy enhances the efficiency of the network without sacrificing performance.
We conduct extensive experiments on benchmark datasets to showcase the effectiveness of our proposed method and their robust generalization across various tasks, including indoor semantic segmentation, outdoor semantic segmentation Our method consistently achieves competitive results compared to existing point transformer-based approaches. Particularly noteworthy is our method’s achievement of state-of-the-art performance, attaining the remarkable mIoU of 73.9% on the S3DIS Area 5 and 62.4% on the SWAN dataset without using any extra data.
@@ II Related Work
II-A Transformer-based Methods
Aiming to maximize the preservation of geometric information in point clouds, state-of-the-art methods prefer to directly process the raw point clouds. The development of the most important unit (i.e., local aggregation) in the point cloud processing network can be broadly divided into three categories, including MLP-, Convolution- and Transformer-based approaches. MLP-based methods [26, 24] employ shared MLPs to extract point-wise features and typically rely on symmetric functions (e.g., max-pooling) to aggregate them. While simple and efficient, they struggle to capture rich local structures that are critical for vision tasks. Motivated by 2D convolution, point convolution methods [34, 18] extend convolutional operations to irregular point sets or point graphs. These approaches dynamically generate convolutional kernels based on local geometric relationships. However, since the weights are learned primarily from low-level geometric cues (e.g., coordinates or relative positions), their ability to capture higher-level contextual dependencies remains limited.
@@ II-A Transformer-based Methods
Unlike convolution-based methods, which learn convolutional weights from low-level point coordinates, attention mechanisms learn attention weights from the connections between point features, thereby exploiting high-level contextual information. Motivated by the success of attention mechanisms in natural language processing and image processing tasks, early methods applied self-attention to global points through scalar dot-product [7, 40], but suffered from high computational costs. These early attention-based methods did not demonstrate superior performance due to the lack of employing position encoding.
@@
The Point Transformer [45] introduces local vector attention to the local points. Additionally, it emphasizes the significance of point position encoding. Subsequent work, Point Transformer V2 [36], incorporates multi-grouping into vector attention inspired by the multi-head strategy. It also enhances position encoding by introducing an additional multiplier to the relation vector, which facilitates learning complex positional relations. To exploit long-range contextual information, the Stratified Transformer [16] densely selects nearby points over a cubic window and sparsely selects distant points. This stratified strategy enlarges the effective receptive field without incurring too much computational overhead. However, by using a window-based approach, the Stratified Transformer focuses on an expanded local region rather than the global region. Moreover, hyper-parameters such as the window size and number of distant points must be optimized for different datasets with varying densities.
@@ II-B Up-sampling in 3D Point Clouds
The hierarchical architecture of a network is instrumental in learning long-range contextual information through down-sampling and grouping operations. Up-sampling involves interpolating new points between known points and adjusting the features of these new points based on their mutual distance to propagate the learned context features to each point. Compared to the extensive research on down-sampling [11, 40] and grouping operations [7], few works specifically emphasize the up-sampling operation. In the Point Transformer [45], the transition up module uses an interpolation operation to recover new point features from known point features through indexing and then integrates these features with those from the encoding layer via a skip connection. Similarly, in Point Transformer V2 [36], the fusion of encoding and decoding layer features is achieved through a skip connection. The distinguishing factor here is that the new point features are unpooled by grid unpooling instead of interpolation. While these methods are simple, they lack semantic awareness and ignore the contextual connection between the encoding and decoding layers.
Our survey highlights several gaps of in current point cloud processing methods. Their fundamental unit, such as point MLP, convolution, transformer, completely ignores task-level information which then leads to sub-optimal performance in downstream tasks. Moreover, the upsampling methods do not effectively communicate between the encoding layers over different resolution points to refine the context information for high-level tasks. Motivated by these gaps, we propose a soft-masked transformer and skip-attention-based upsampling. Moreover, we propose shared point position encoding to reduce the network parameters and training time.
@@ III Method
We revisit the two classical local vector attention-based point transformers, namely Point Transformer and Point Transformer V2, in Section III-A. Then, we present the SMTransformer block in Section III-B, followed by our skip attention-based up-sampling block in Section III-C. We introduce the shared point position encoding strategy in Section III-D, and finally, in Section III-E, we provide details about our network.
@@ III-A Rethinking Vector Attention based Transformer
$\mathcal{G}_{i}=\sum_{j=1}^{K}\mathcal{A}\big((K_{ij}^{f}\ominus Q_{i}^{f})\oplus\delta_{b}\big)\odot\big(V_{ij}^{f}\oplus\delta_{b}\big),\vskip-5.69054pt$ (1)
where $Q_{i}^{f}\in\mathbb{R}^{{N\times C}}$ is the query matrices of $f_{i}$. $K_{ij}^{f}$ and $V_{ij}^{f}\in\mathbb{R}^{{N\times K\times C}}$ are the key and value matrices of $f_{ij}$. The subtraction operation $\ominus$ between $Q_{i}^{f}$ and $K_{ij}^{f}$ is performed via broadcasting to ensure dimension compatibility. $\delta_{b}=\delta_{b}(\Delta p_{ij})$ is the position encoding bias. $\mathcal{A}(\cdot)$ denotes the local vector attention function, implemented by MLP, followed by the softmax. $\sum$ stands for the symmetric operation (SOP) (e.g. summation) and $\odot$ is the element-wise multiplication operation. $\oplus$ is the element-wise addition operation and $\mathcal{G}_{i}$ is the output feature of transformer.
@@
$\mathcal{G}_{i}=\sum_{j=1}^{K}\mathcal{A}\big({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{m}}(K_{ij}^{f}\ominus Q_{i}^{f})\oplus\delta_{b}\big)\odot\big(V_{ij}^{f}\oplus\delta_{b}\big),\vskip-5.69054pt$ (2)
where $\delta_{m}=\delta_{m}(\Delta p_{ij})$ is the position encoding multiplier. The attention function $\mathcal{A}(\cdot)$ learns robust weights from the rich relationships, including the low-level geometric relationship (i.e. relation position) and the high-level contextual relationship (i.e. relation feature). Due to the inherent unstructured nature of point clouds, local neighborhoods are typically constructed based on purely position and feature, However, this approach is agnostic to semantic boundaries and often leads to the incorrect aggregation of points from semantically distinct regions.
@@ III-B Soft Masked Transformer Block (SMTB)
To address above problem arising from the point cloud data structure, we designed the soft mask transformer. By introducing a learnable semantic signal, it dynamically adjusts attention weights within these geometric neighborhoods, enabling the network to perceive and respect semantic boundaries. The soft-masked transformer can be expressed as,
@@ III-B1 Soft Mask
where the $W_{q}^{s}$ and $W_{k}^{s}$ ($\mathbb{R}^{N\times C}\rightarrow\mathbb{R}^{N\times T}$) are learnable weight matrices of a shared linear layer that projects the $C$-dimensional feature of each point into a $T$-dimensional task score vector, where $T$ is the number of the classes. These matrices are implemented as 1x1 convolution layers with weight shapes of (C, T) and are optimized end-to-end via backpropagation. The output is passed through a Softmax function along the T-dimension to produce class probabilities. $G[\cdot]$ is the grouping operation to obtain the task scores of neighbour points at point $p_{i}$. The soft mask is generated from the score difference $(K_{ij}^{s}\ominus Q_{i}^{s})\in\mathbb{R}^{{N\times K\times T}}$ as,
[…]
where the ${\rm Max}(\cdot)$ is the maximum, ${\rm Norm}(\cdot)$ is Min-Max Normalization, and $||\cdot||_{2}$ is the Euclidean Norm.
The resulting soft mask $S(\cdot)$ modulates the original attention weights. A high mask value indicates a strong semantic boundary and is used to down-weight attention to that neighbor, preventing feature corruption from adjacent but semantically distinct regions. This enhances robustness at class boundaries. The module’s design is inherently general, as it relies on a learnable task signal (class scores), not on dataset-specific geometry. It can serve as a general-purpose component for point cloud processing.